# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v2

@@ 3.1 Erdős-Rényi model
In this subsection, we focus on the Erdős-Rényi model and use $\hat{\pi}$ from (6) with $f(x,y)=xy$ as the estimator. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])\sim\mathrm{Bin}(N_{k},p_{11})$ with $|F|=k$, the error probability follows from the standard Chernoff bound (see, e.g., (52)): for $10<\eta<$ 1,
@@
For the bad event of noise, applying (8) along with the lower-order cumulant (36) and (37) yields the following lemma, whose proof is deferred to Appendix B.2.
@@
Since $p_{11}=p^{2}(1+\gamma)$, we need to choose an appropriate $\eta\in(0,1)$ depending on the model parameters, such that $(1+\gamma)(1-\eta)>1$, in order to apply the error probabilities in (13) and (14) to upper bound (11) and (12), respectively. When $m$ exceeds the sample complexities for partial and exact recovery, we explicitly construct $\eta$ to prove the following propositions, which establish the possibility results stated in Theorems 1 and 2. The proofs of Propositions 1 and 2 are deferred to Appendix A.1 and A.2, respectively.
@@ 3.2 Gaussian Wigner model
In this subsection, we focus on the Gaussian Wigner model. To establish upper bounds for exact recovery, we use different estimators depending on the value of $\rho$, where $0<\rho\leq 1-e^{-12}$ corresponds to the weak signal regime, and $11-e^{-12}<\rho<$ 1 corresponds to the strong signal regime.
@@
the tail follows from the Hanson-Wright inequality in Lemma 9. For the bad event of noise, applying (8) with the formula of lower-order cumulants (41) and (42) yields the following Lemma, whose proof is deferred to Appendix B.3.
@@
The following proposition provides sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $0<\rho\leq 1-e^{-12}$, whose proof is deferred to Appendix A.3.
@@
Under the strong signal regime $11-e^{-12}<\rho<$ 1, we use $\hat{\pi}$ in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ as our estimator. Indeed, the MLE is also of the form (6), but with $f(x,y)=\rho xy-\frac{\rho^{2}}{2}(x^{2}+y^{2})$, which is approximately $-\frac{1}{2}(x-y)^{2}$ when $\rho=1-o(1)$. This choice reflects the phase transition phenomenon in the Gaussian Wigner model when $\rho=1-o(1)$. Specifically, for a correlated edge pair, the expected mean squared difference $\mathbb{E}\left[\left(\beta_{e}(G_{1})-\beta_{\pi^{*}(e)}(G_{2})\right)^{2}\right]=o(1)$, while for an uncorrelated edge pair, it stays bounded away from zero with high probability. Our proof follows a similar structure with the weak signal regime. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])=\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}$, where $(A_{i},B_{i})$ are i.i.d. pairs of standard normals with correlation coefficient $\rho$. Then, the error probability can be upper bounded by
@@
the tail follows from the concentration inequality for chi-squared distribution in Lemma 10. For the bad event of noise, applying (8) with the formula of lower-order cumulants (45) and (46) yields the following Lemma, whose proof is deferred to Appendix B.4.
@@
The following proposition provides a sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $11-e^{-12}<\rho<$ 1, whose proof is deferred to Appendix A.4.
@@
When $11-e^{-12}<\rho<$ 1, there exists a universal constant $C_{4}>0$ such that, when $m\geq C_{4}\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ satisfies
@@
In view of Propositions 3 and 4, we note that $\rho^{2}\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $0<\rho\leq 1-e^{-12}$, and $\log\left(\frac{1}{1-\rho}\right)\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $11-e^{-12}<\rho<$ 1. Then, there exists a universal constant $C_{2}$ such that,
@@
Our estimator (6) requires knowledge of the cardinality $|S^{*}|=m$. When $m$ is unknown, the objective function in (6) will increase monotonically with $m$. To address this, we can introduce a penalty term for $m$ in the estimator to identify the correct size, following an approach similar to the MLE in the new model. This will be discussed in more detail in a forthcoming working paper.
@@ 4 Impossibility results
In this section, we present the impossibility results for the graph alignment problem. Under our proposed model, the alignment problem aims to recover the domain $S^{*}\subseteq V(G_{1})$, range $T^{*}\subseteq V(G_{2})$, and the mapping $\pi^{*}:S^{*}\mapsto T^{*}$. When equipped with the additional knowledge on $S^{*}$ and $T^{*}$, our problem can be reduced to recovery with full observations on smaller graphs, the reconstruction threshold for which is settled in [WXY22]. The lower bound therein remains valid when the number of correlated nodes is substituted with $m$. However, such reduction only proves tight in a limited number of regimes (see Proposition 6). We will establish the impossibility results for the remaining regimes by Fano’s method. Two main ingredients of Fano’s method are outlined as follows:
@@
  • Construct a packing set ${\mathcal{M}}$ of the parameter space $\mathcal{S}_{n,m}$ such that any two distinct elements from ${\mathcal{M}}$ differ by a prescribed threshold. Specifically, in partial recovery, the overlap of each pair is less than $\delta$, which is equivalent to $\min_{\pi\neq\pi^{\prime}\in{\mathcal{M}}}d(\pi,\pi^{\prime})\geq(1-\delta)m$, while in exact recovery ${\mathcal{M}}=\mathcal{S}_{n,m}$. The cardinality of ${\mathcal{M}}$ measures the complexity of the parameter space under the target metric.
@@
  • Choose the uniform prior on $\pi^{*}$ over ${\mathcal{M}}$ and upper bound the mutual information $I(\pi^{*};G_{1},G_{2})$. Given $\pi^{*}$, the conditional distribution of the observed graphs $(G_{1},G_{2})$ is specified in Definition 4. For the mutual information, let $\mathcal{P}$ denote the joint distribution of $(G_{1},G_{2})$ and $\mathcal{Q}$ be any distribution over $(G_{1},G_{2})$. Then,
    $I(\pi^{*};G_{1},G_{2})=\mathbb{E}_{\pi^{*}}[D(\mathcal{P}_{G_{1},G_{2}|\pi^{*}}\|\mathcal{P}_{G_{1},G_{2}})]\leq\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}).$ (18)
    The impossibility results follow if $I(\pi^{*};G_{1},G_{2})\leq c\log|{\mathcal{M}}|$ for some small constant $c$.
Let ${\mathcal{M}}_{\delta}$ denote a packing set under the overlap threshold $\delta$. The size of ${\mathcal{M}}_{\delta}$ follows from the standard volume argument [PW25, Theorem 27.3]. For $r\in[m]$, let $B(\pi,r)\triangleq\{\pi^{\prime}:d(\pi,\pi^{\prime})\leq r\}$ denote the ball of radius $r$ centered at $\pi$. Then, we obtain
It remains to evaluate the cardinality of $\mathcal{S}_{n,m}$ and upper bound the volume of the ball under our distance metric $d$. It is straightforward to obtain that $|\mathcal{S}_{n,m}|=\binom{n}{m}^{2}m!.$ Let $k=\delta m$. Note that all elements from $B(\pi,m-k)$ have at least $k$ common mappings. To upper bound $|B(\pi,m-k)|$, we first choose $k$ elements from the domain of $\pi$ and map to the same value as $\pi$, and the remaining domain and range of size $m-k$ and the mapping are selected arbitrarily. We get $|B(\pi,m-k)|\leq\binom{m}{k}\binom{n-k}{m-k}^{2}(m-k)!.$ Consequently,
@@
where we use the inequalities that $(\frac{n}{k})^{k}\leq\binom{n}{k}<(\frac{en}{k})^{k}$ and $k!\geq(k/e)^{k}$. Fano’s method provides a lower bound on the Bayesian risk when $\pi$ is uniformly distributed over ${\mathcal{M}}_{\delta}$, which further lower bounds the minimax risk. The above argument also yields a lower bound when $\pi$ is uniform over $\mathcal{S}_{n,m}$ via generalized Fano’s inequality [BHM12, Lemma 20]. For the Erdős-Rényi model, the following propositions provide lower bounds for $m$ for partial recovery and exact recovery, and thus prove the lower bounds in Theorems 1 and 2. For the Gaussian Wigner model, the Proposition 5 provides a lower bound for $m$ for partial recovery, which implies that $m\leq\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}$ is sufficient for the lower bound since $m<1$ is impossible for partial recovery. Consequently, we finish the proof of impossibility results in Theorem 3 since $\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}\asymp\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$.
@@
where $P=\mathrm{Bern}(p,p,\rho)$ and $Q=\mathrm{Bern}(p,p,0)$. Pick $\mathcal{Q}$ in (18) to be an auxiliary null model under which $G_{1}$ and $G_{2}$ are independent with the same marginal as $\mathcal{P}$. Then, $\mathcal{Q}_{G_{1},G_{2}}$ can be factorized into
for any $\pi:S\mapsto T$ with $|S|=|T|=m$. Applying Lemma 7, we obtain
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq 25\binom{m}{2}p^{2}\phi(\gamma).$ (20)
Applying generalized Fano’s inequality [BHM12, Lemma 20] with (19) and (20), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{25\binom{m}{2}p^{2}\phi(\gamma)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{13c}{\delta},$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
Applying Lemma 7, we obtain
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$ (21)
Applying generalized Fano’s inequality [BHM12, Lemma 20] with (19) and (21), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{c}{2\delta},$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$. ∎
@@
For any $c\in(0,1)$ and any estimator $\hat{\pi}$, there exists a constant $c_{3}$ depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)$,
@@
When $\gamma\geq 1$, $\gamma\wedge\gamma^{2}=\gamma$. By comparing (22) and (23), we derive that exact recovery is impossible if $m\lesssim\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)$. ∎

### v3

@@ 3.1 Erdős-Rényi model
In this subsection, we focus on the Erdős-Rényi model and use $\hat{\pi}$ from (6) with $f(x,y)=xy$ as the estimator. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])\sim\mathrm{Bin}(N_{k},p_{11})$ with $|F|=k$, the error probability follows from the standard Chernoff bound (see, e.g., (57)): for $0<\eta<1$,
@@
For the bad event of noise, applying (8) along with the lower-order cumulant (38) and (39) yields the following lemma, whose proof is deferred to Appendix C.2.
@@
Since $p_{11}=p^{2}(1+\gamma)$, we need to choose an appropriate $\eta\in(0,1)$ depending on the model parameters, such that $(1+\gamma)(1-\eta)>1$, in order to apply the error probabilities in (14) and (15) to upper bound (12) and (13), respectively. When $m$ exceeds the sample complexities for partial and exact recovery, we explicitly construct $\eta$ to prove the following propositions, which establish the possibility results stated in Theorems 1 and 2. The proofs of Propositions 1 and 2 are deferred to Appendix B.1 and B.2, respectively.
@@ 3.2 Gaussian Wigner model
In this subsection, we focus on the Gaussian Wigner model. Different from Section 3.1, the Gaussian Wigner model requires a different estimator as $\rho\to 1$. Intuitively, when $\rho=1$, the edge weights satisfy $\beta_{e}(G_{1})=\beta_{\pi^{*}(e)}(G_{2})\in\mathbb{R}$ for all $e\in\binom{S^{*}}{2}$. Since each weighted edge is marginally standard normal, the ground truth permutation $\pi^{*}$ can be exactly recovered by comparing edge weights pairwise, as long as $m\geq 3$ (the node correspondence is not identifiable when $m=2$). To establish upper bounds for exact recovery, we use different estimators depending on the value of $\rho$, where $0<\rho\leq 1-e^{-12}$ corresponds to the weak signal regime, and $1-e^{-12}<\rho<1$ corresponds to the strong signal regime.
@@
the tail follows from the Hanson-Wright inequality in Lemma 10. For the bad event of noise, applying (8) with the formula of lower-order cumulants (43) and (44) yields the following Lemma, whose proof is deferred to Appendix C.3.
@@
The following proposition provides sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $0<\rho\leq 1-e^{-12}$, whose proof is deferred to Appendix B.3.
@@
Under the strong signal regime $1-e^{-12}<\rho<1$, we use $\hat{\pi}$ in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ as our estimator. Indeed, the MLE is also of the form (6), but with $f(x,y)=\rho xy-\frac{\rho^{2}}{2}(x^{2}+y^{2})$, which is approximately $\frac{1}{2}(x-y)^{2}$ when $\rho=1-o(1)$. This choice reflects the phase transition phenomenon in the Gaussian Wigner model when $\rho=1-o(1)$. Specifically, for a correlated edge pair, the expected mean squared difference $\mathbb{E}\left[\left(\beta_{e}(G_{1})-\beta_{\pi^{*}(e)}(G_{2})\right)^{2}\right]=o(1)$, while for an uncorrelated edge pair, it stays bounded away from zero with high probability. Our proof follows a similar structure with the weak signal regime. For the bad event of signal, since $\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f})-\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F])=\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}$, where $(A_{i},B_{i})$ are i.i.d. pairs of standard normals with correlation coefficient $\rho$. Then, the error probability can be upper bounded by
@@
the tail follows from the concentration inequality for chi-squared distribution in Lemma 11. For the bad event of noise, applying (8) with the formula of lower-order cumulants (47) and (48) yields the following Lemma, whose proof is deferred to Appendix C.4.
@@
The following proposition provides a sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $1-e^{-12}<\rho<1$, whose proof is deferred to Appendix B.4.
@@
When $1-e^{-12}<\rho<1$, there exists a universal constant $C_{4}>0$ such that, when $m\geq C_{4}\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ satisfies
@@
In view of Propositions 3 and 4, we note that $\rho^{2}\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $0<\rho\leq 1-e^{-12}$, and $\log\left(\frac{1}{1-\rho}\right)\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $1-e^{-12}<\rho<1$. Then, there exists a universal constant $C_{2}$ such that,
@@
Our estimator (6) requires knowledge of the cardinality $|S^{*}|=m$. For certain problems, such as induced subgraph sampling, the random variable $m$ can be shown to be highly concentrated by analyzing its hypergeometric distribution [HY25]. When $m$ is unknown, the objective function in (6) will increase monotonically with $m$. To address this, we can introduce a penalty term for $m$ in the estimator to identify the correct size. Specifically, when $m$ is unknown, the estimator takes the form
$\hat{\pi}\in\mathop{\text{argmax}}\limits_{\pi\in\mathcal{S}_{n,m},m\leq n}\left[\mathsf{{e}}(\mathcal{H}_{\pi}^{f})-F(m)\right],$
where $F(m)=\lambda m^{2}$ is a penalty term and $\lambda$ is a tuning parameter. Indeed, we choose $F(m)$ such that the estimator encourages correlated pairs and penalizes independent pairs in order to correctly identify $m$.
@@ 4 Impossibility results
In this section, we present the impossibility results for the graph alignment problem. Under our proposed model, the alignment problem aims to recover the domain $S^{*}\subseteq V(G_{1})$, range $T^{*}\subseteq V(G_{2})$, and the mapping $\pi^{*}:S^{*}\mapsto T^{*}$. When equipped with the additional knowledge on $S^{*}$ and $T^{*}$, our problem can be reduced to recovery with full observations on smaller graphs, the reconstruction threshold for which is settled in [WXY22]. The lower bound therein remains valid when the number of correlated nodes is substituted with $m$. However, such reduction only proves tight in a limited number of regimes (see Proposition 6). We will establish the impossibility results for the remaining regimes by Fano’s method (see, e.g., [Fan61],  [Yu97], and [CT06, Section 2.10]), which consists of the following two ingredients:
  • Separation. Construct a subset of well-separated parameters. Any testing error then leads to an estimation error proportional to the minimum separation between parameters.
  • Insufficient information. The mutual information between the underlying parameter and the observed data is small. In this case, any test incurs a non-negligible error probability.
The technical outlines of the Fano’s inequality are sketched below.
Constructing a subset of parameters.
@@
Let ${\mathcal{M}}_{\delta}$ be a packing set of $\mathcal{S}_{n,m}$ such that any two distinct elements $\pi,\pi^{\prime}\in{\mathcal{M}}$ differ by a prescribed threshold. Specifically, in partial recovery, we need $\min_{\pi\neq\pi^{\prime}\in{\mathcal{M}}}d(\pi,\pi^{\prime})>(1-\delta)m$; in exact recovery, we simply choose ${\mathcal{M}}_{1}=\mathcal{S}_{n,m}$. The cardinality of ${\mathcal{M}}_{\delta}$ measures the complexity of the parameter space under the target metric.
Bounding the mutual information $I(\pi^{*};G_{1},G_{2})$.
@@
Given $\pi^{*}$, the conditional distribution of the observed graphs $(G_{1},G_{2})$ is specified in Definitions 4 and 5. For the mutual information, let $\mathcal{P}$ denote the joint distribution of $(G_{1},G_{2})$ and $\mathcal{Q}$ be any distribution over $(G_{1},G_{2})$. Then,
$I(\pi^{*};G_{1},G_{2})=\mathbb{E}_{\pi^{*}}[D(\mathcal{P}_{G_{1},G_{2}|\pi^{*}}\|\mathcal{P}_{G_{1},G_{2}})]\leq\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}),$ (19)
where the inequality is because
$D(\mathcal{P}_{G_{1},G_{2}|\pi^{*}}\|\mathcal{P}_{G_{1},G_{2}})=D(\mathcal{P}_{G_{1},G_{2}|\pi^{*}}\|\mathcal{Q}_{G_{1},G_{2}})-D(\mathcal{P}_{G_{1},G_{2}}\|\mathcal{Q}_{G_{1},G_{2}})$
and the KL divergence $D(\mathcal{P}_{G_{1},G_{2}}\|\mathcal{Q}_{G_{1},G_{2}})\geq 0$ for any distribution $\mathcal{Q}$.
Applying Fano’s inequality.
By Fano’s inequality, with $\pi^{*}$ being the discrete uniform prior in the packing set $\mathcal{M}_{\delta}$, for any estimator $\hat{\pi}$, we have
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\geq 1-\frac{I(\pi^{*};G_{1},G_{2})+\log 2}{\log|\mathcal{M}_{\delta}|},$
where $I(\pi^{*};G_{1},G_{2})$ denotes the mutual information between $\pi^{*}$ and $(G_{1},G_{2})$, and $\pi^{*}$ is uniformly distributed over $\mathcal{M}_{\delta}$.
The relevant quantities are evaluated in the next lemma.
Lemma 5.
For any $0<\delta<1$, we have
$|\mathcal{M}_{\delta}|\geq\left(\frac{\delta n}{e^{3}}\right)^{\delta m}.$ (20)
In the Erdős-Rényi model, we have
$I(\pi^{*};G_{1},G_{2})\leq 25\binom{m}{2}p^{2}\phi(\gamma).$ (21)
In the Gaussian Wigner model, we have
$I(\pi^{*};G_{1},G_{2})\leq\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$ (22)
@@
The proof of Lemma 5 is deferred to Appendix C.5. Fano’s method provides a lower bound on the Bayesian risk when $\pi$ is uniformly distributed over ${\mathcal{M}}_{\delta}$, which further lower bounds the minimax risk. The above argument also yields a lower bound when $\pi$ is uniform over $\mathcal{S}_{n,m}$ via Fano’s inequality. For the Erdős-Rényi model, the following propositions provide lower bounds for $m$ for partial recovery and exact recovery, and thus prove the lower bounds in Theorems 1 and 2. For the Gaussian Wigner model, the Proposition 5 provides a lower bound for $m$ for partial recovery, which implies that $m\leq\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}$ is sufficient for the lower bound since $m<1$ is impossible for partial recovery. Consequently, we finish the proof of impossibility results in Theorem 3 since $\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}\asymp\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$.
@@
Applying Fano’s inequality with (20) and (21), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{25\binom{m}{2}p^{2}\phi(\gamma)+\log 2}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{13c}{\delta},$
where $\pi^{*}$ is uniformly distributed over $\mathcal{M}_{\delta}$.
Applying Fano’s inequality with (20) and (22), we obtain
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right)+\log 2}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{c}{2\delta},$
where $\pi^{*}$ is uniformly distributed over $\mathcal{M}_{\delta}$. ∎
@@
For any $c\in(0,1)$ and any estimator $\hat{\pi}$, there exists a constant $c_{3}$ depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\left(\frac{1}{p^{2}\gamma}\log\frac{1}{p^{2}\gamma}\right)\right)$,
@@
When $\gamma\geq 1$, $\gamma\wedge\gamma^{2}=\gamma$. By comparing (23) and (24), we derive that exact recovery is impossible if $m\lesssim\frac{\log n}{p^{2}\phi(\gamma)}\vee\left(\frac{1}{p^{2}\gamma}\log\frac{1}{p^{2}\gamma}\right)$. ∎
Remark 6.
As discussed in Section 1.2, the estimation of $\pi^{*}$ consists of two different subproblems: 1) recovering the support sets $S^{*}$ and $T^{*}$; and 2) recovering the mapping $\pi^{*}$ between these two sets. The first lower bound $\frac{\log n}{p^{2}\phi(\gamma)}$ arises from the overlap complexity associated with the former (see Appendix A for a proof), while the second lower bound $\frac{1}{p^{2}\gamma}\log\frac{1}{p^{2}\gamma}$ is due to the latter.